• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kumbartzki

    Log entry time 15:56:13 on February 1,2002

    Entry number 81485

    keyword=day shift summary

    Day shift summary:
    Continued data taking f* Runs 1950-1952
    10:10 Access to hall to install post radiator.
    11:35 resummed data taking
    12:30 changed to Carbon target for optics run.
    12:50 finished running, Access to hall for polarity change on HRS,
    (including repairs to cables), Calo moved.
    Changed spectormeter angel to 14.57 and momentum setting to 4.0 GeV.
    15:21 resumed data taking on C tgt with ~50 uA.